(D+3)^2y=sinh2x


1
Expert's answer
2021-06-01T12:22:07-0400

Ans:-"(D+3)^2y=Sinh2x"

Auxiliary equation

"(D+3)^2=0\\"

"D=-3,-3" are the roots

"\therefore C.F.=(C_1+C_2x)e^{-3x}"


"Sinhx=\dfrac{e^x-e^{-x}}{2}"

Particular Integral


"P.I.=\dfrac{Sinh2x}{(D+3)^2}"


"=\dfrac{1}{2}[\dfrac{e^{2x}-e^{-2x}}{(D+3)^2}]"


"=\dfrac{1}{2}\dfrac{e^{2x}}{(D+3)^2}-\dfrac{1}{2}\dfrac{e^{-2x}}{(D+3)^2}"


"=\dfrac{1}{2}\dfrac{e^{2x}}{(2+3)^2}-\dfrac{1}{2}\dfrac{e^{-2x}}{(-2+3)^2}"


"=\dfrac{e^{2x}}{50}-\dfrac{e^{-2x}}{2}"


Hence the complete solution

"Y=C.F.+P.I"


"Y=(C_1+C_2x)e^{-3x}+\ \dfrac{e^{2x}}{50}-\dfrac{e^{-2x}}{2}"